OpenStudy (anonymous):

Find the discriminant of the polynomial below. 9x^2-18x+9

OpenStudy (anonymous):

648

OpenStudy (anonymous):

\[-18^{2}-4(9)(9)=-324-324=648\]

OpenStudy (anonymous):

\[take \left( -18 \right)^{2}=324\]

OpenStudy (anonymous):

so 0

OpenStudy (anonymous):

so it is 324-324=0

OpenStudy (mathstudent55):

\(b^2 - 4ac =\) \(= (-18)^2 - 4( 9)(9) \) \(= 324 - 324 \) \(= 0\)

OpenStudy (mathstudent55):

You need to be careful with this: If a = 5, then you know that a^2 = 5^2 = 25. That's simple. Here's where you need to be careful. If a = -5, then a^2 = (-5)^2 = 25. You must have the parentheses around -5. This is correct. The following line is incorrect: If a = -5, then a^2 = -5^2 = -25. In math, (-5)^2 means (-5)(-5) = 25. In math, -5^2 means -(5)(5) = -25. If you have a = -5, and you need to find a^2, you need to do (-5)^2 which means (-5)(-5) which equals positive 25, since it is the product of 2 negative numbers. If you incorrectly state that if a = -5, then a^2 = -5^2, then you'll get -25 which is incorrect.